What is the present worth difference between an investment of $35,000 per year for 50 years and an investment of $35,000 per year forever at an interest rate of 9% per year?

The difference is determined to be $_______.
Answer #1

PW of the investment for 50 years = 35000(P/A,9%,50) = 35000(10.96168) = 383658.8

PW of the investment per year forever =35000/0.09 = 388888.88

Difference = 388888.88-383658.8 = 5230.08